OpenStudy (anonymous):

The following table shows the length and width of two rectangles:

OpenStudy (anonymous):

Length Width Rectangle A 3x + 2 2x - 1 Rectangle B x + 5 5x - 1

OpenStudy (anonymous):

Which expression is the result of the perimeter of rectangle B minus the perimeter of rectangle A?

OpenStudy (anonymous):

2x + 2 2x + 6 22x + 14 22x + 18

OpenStudy (anonymous):

Perimeter= 2(length)+2(width)

OpenStudy (anonymous):

I know how to do it, I'm just trying to get this assignment done in a short amount of time

OpenStudy (uri):

First you need to take the perimeter of the rectangle,do you know how to? P=2l+2w That would be P=2( x + 5)+2(5x - 1) Solve it and do this with Rectangle A as well.

OpenStudy (anonymous):

So we'd have P(B)-P(A) where P(B)= the perimeter of B and P(A)= the perimeter of A \[\Large (2(x+5)+2(5x-1))-(2(3x+2)+2(2x-1))\]

OpenStudy (anonymous):

length times width=Area, not perimeter

OpenStudy (uri):

@SiennaSurfer Well you have to do it yourself. We are here to help you..not do your work.... Please : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

oh... right, thx @doulikepiecauseidont

OpenStudy (anonymous):

and @uri , I know, I was just trying to make it a little faster

OpenStudy (uri):

Well It's against the Coc. Next time be careful.

OpenStudy (anonymous):

alright. thx ;)

OpenStudy (anonymous):

Wait so what was the answer?